# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ B.1 Moment ODEs and analytic solutions
(iii). Note in the below, with a slight abuse of notation, we write $A_{n}$ to denote the value of the matrix $A$ at the $n$th iteration and $A_{ij}$ for the $i$th row and $j$th column element of the matrix $A$. We will demonstrate the reasoning for $\mathrm{KL}(\mu_{n\gamma}||\pi)$, noting that the same calculations hold for $\mathrm{KL}(\nu_{n}||\pi)$ and $\mathrm{KL}(\eta_{n}||\pi)$ due to Lemma 3.1 with $E_{n}$ replaced by $E_{n}^{\beta}$ and $E_{n}^{\alpha}$ respectively.
Let $A_{n}:=C_{\pi}^{-1}E_{n}$ and using Lemma 3.1 and (B.24) we obtain
$\mathrm{KL}(\mu_{n\gamma}||\pi)$ $=\frac{1}{2}\left(-\log\det(I+A_{n})+Tr[A_{n}]+Tr[A_{0}^{-1}C_{\pi}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}A_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}A_{n}^{2}e^{n\gamma C_{\pi}^{-1}}]\right)$
$=\frac{1}{2}\left(-\log\det(I+A_{n})+Tr[A_{n}]+Tr[F_{n}A_{n}^{2}]\right)$
where $F_{n}:=e^{n\gamma C_{\pi}^{-1}}D_{0}e^{n\gamma C_{\pi}^{-1}}$ and $D_{0}:=A_{0}^{-1}C_{\pi}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}A_{0}^{-1}$ and recall that using (3.10)
$A_{n}$ $=e^{-n\gamma\Gamma}\left(A_{0}^{-1}+\Omega\,(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$
@@
Denote $K_{n}:=\left(A_{0}^{-1}+\Omega\,(I-e^{-2n\gamma\Gamma})\right)^{-1}$ so that $A_{n}=e^{-n\gamma\Gamma}K_{n}e^{-n\gamma\Gamma}$. Since $A_{n}=C_{\pi}^{-1}C_{n\gamma}-I$ has eigenvalues larger than 1-1, we have $\det(I+A_{n})=e^{Tr[\log(I+A_{n})]}$, and
$-\log\det(I+A_{n})=-Tr[\log(I+A_{n})]=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$
@@
whenever $A_{n}||<1$ (so that the series converges), which holds for sufficiently large $n$. Then
$2\mathrm{KL}(\mu_{n\gamma}||\pi)=Tr[F_{n}A_{n}^{2}]+\frac{1}{2}Tr[A_{n}^{2}]+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$
where $\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$ is convergent and equal to $\log\det(I+A_{n})-Tr[A_{n}^{2}]/2+Tr[A_{n}]$. Also, $\Gamma$ is diagonalisable when $C_{\pi}$ is positive definite and we denote its spectral decomposition by $\Gamma=P^{\top}\Lambda P$ with $\lambda_{i}=\Lambda_{ii}$ denoting the $i$th eigenvalue of $\Gamma$ and $\frac{1}{2}<\lambda_{1}<\lambda_{2}<\lambda_{3}\dots<\lambda_{d}$ and $p_{i}$ the $i$th column of $P$ where $P$ is unitary. Then
$Tr[A_{n}^{2}]$ $=Tr[(P^{\top}e^{-n\gamma\Lambda}PK_{n}P^{\top}e^{-n\gamma\Lambda}P)^{2}]$
$=Tr[e^{-2n\gamma\Lambda}Me^{-2n\gamma\Lambda}M]$
$=\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}M_{ji}$
where $M:=PK_{n}P^{\top}$ since $Tr[P^{\top}DP]=\sum_{i=1}^{d}D_{ii}$ for $P$ unitary orthonormal matrix, $PP^{\top}=I$. A similar representation can be obtained for $Tr[A_{n}^{k}]$ involving $k$ iterated sums. Similarly,
$Tr[F_{n}A_{n}^{2}]$ $=e^{-n\gamma}Tr[K_{n}D_{0}K_{n}e^{-2n\gamma\Gamma}]=e^{-n\gamma}Tr[e^{-2n\gamma\Lambda}PK_{n}D_{0}K_{n}P^{\top}]=\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}$
where $N:=PK_{n}D_{0}K_{n}P^{\top}$. Using the same reasoning for $\mathrm{KL}(\nu_{n}||\pi)$ yields
$\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ $=\frac{\frac{1}{2}\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}^{\beta}M_{ji}^{\beta}+\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}^{\beta}+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[B_{n}^{k}]}{\frac{1}{2}\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}M_{ji}+\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]}$
$=\frac{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}^{\beta}+\mathcal{O}(e^{-4n\gamma\lambda_{1}})}{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}+\mathcal{O}(e^{-4n\gamma\lambda_{1}})}$
$=\frac{N_{11}^{\beta}+\mathcal{O}(e^{-n\gamma\min(4\lambda_{1},2(\lambda_{2}-\lambda_{1}))})}{N_{11}+\mathcal{O}(e^{-n\gamma\min(4\lambda_{1},2(\lambda_{2}-\lambda_{1}))})}$
@@
where $M^{\beta}=PK_{n}^{\beta}P^{\top}$, $N^{\beta}:=PK_{n}^{\beta}D_{0}K_{n}^{\beta}P^{\top}$, $K_{n}^{\beta}:=(A_{0}^{-1}+\Omega^{\beta}(I-e^{-2n\gamma\Gamma}))^{-1}$ and $B_{n}=e^{-n\gamma\Gamma}K_{n}^{\beta}e^{-n\gamma\Gamma}$ and in the third line we have divided top and bottom by $e^{-2n\gamma(\lambda_{1}+\frac{1}{2})}$ and noting that $K_{n}\rightarrow(A_{0}^{-1}+\Omega)^{-1}$ and $K_{n}^{\beta}\rightarrow(A_{0}^{-1}+\Omega^{\beta})^{-1}$ as $n\rightarrow\infty$. Taking the limit of the above expression as $n\rightarrow\infty$ yields the result.
@@ Appendix C Preservation of log-concavity
Our proof strategy involves starting with a sequential splitting of the WFR flow and using Girsanov theorem to characterise the intermediate density due to the W flow. This is based on a similar proof strategy used to show that the filtering density satisfies a Poincaré inequality (see Lemma 5.1 in [Pathiraja2021]). In Appendix C.1, we show that when the initial density is “sufficiently” strongly log-concave (to be made precise), there exists a finite time interval over which the W flow maintains log-concavity. Although the W flow cannot be expected to preserve log-concavity uniformly in time (unless $\pi$ is Gaussian), the properties of the FR flow can be exploited to maintain log-concavity uniformly in time. The next lemma shows that the FR does exactly this, and then the proof in Appendix C.2 combines the results.
@@
Proof.
The first step will be to use Girsanov theorem to characterise the law of the overdamped Langevin diffusion, using the re-scaled Langevin diffusion
$dX_{t}=\nabla U(X_{t})dt+dW_{t}$
where $U=-\frac{1}{2}V_{\pi}$. Let $\widehat{\mu}_{t}$ denote the law of $X_{t}$, so then
$\widehat{\mu}_{t}$ $=\int_{C[0,t]}\exp\left(\int_{0}^{t}\nabla U(W_{s})dWs-\frac{1}{2}\int_{0}^{t}|\nabla U(W_{s})|^{2}ds\right)\mu_{W}(dw)$
$=\int_{C[0,t]}\exp\left((U(W_{t})-U(W_{0})-\frac{1}{2}\int_{0}^{t}(\Delta U+|\nabla U|^{2})(W_{s})ds\right)\mu_{W}(dw)$
where $\mu_{W}(dw)$ denotes the Wiener measure on path space $C[0,t]$ and in the second line we have used Itô formula. Here the application of Girsanov theorem is valid since Novikov’s condition is satisfied under Assumption 1. Consider now a discretisation of the time interval $[0,t]$ with step size $\tau$ and $N$ steps, $0=t_{0},t_{1},t_{2},\dots,t_{N-1},t_{N}=t$ with $t_{i}-t_{i-1}=\tau$ for all $i=1,2,3,\dots,N$. From now on, we use the notation $\widehat{\mu}_{i}$ to denote an approximation to $\widehat{\mu}_{t_{i}}$ (and likewise for other quantities). We can then approximate as
$\widehat{\mu}_{t}\approx$ $\int_{C[0,t]}\exp\left(\frac{1}{2}(V_{\pi}(w_{0})-V_{\pi}(w_{N}))\right)\prod_{i=1}^{N-1}\exp\left(-\frac{1}{2}R(w_{i})\right)\mu_{W}(dw)$
@@
$\mu_{i}(w_{i}):=\left\{\begin{array}[]{lr}\int Q(w_{{i-1}})\,q_{\tau}(w_{{i-1}},w_{i})\,\mu_{i-1}(w_{i-1})\,dw_{{i-1}},&i=2,3,\dots\\ \int Q(w_{0})\,q_{\tau}(w_{0},w_{i})\,\exp\left(-U(w_{0})\right)\,\exp(-V_{0}(w_{0}))\,dw_{0},&i=1\end{array}\right$.
@@
$R(w)$ $:=\Delta U(w)+|\nabla U(w)|^{2}$
@@
$=\exp(-\mathcal{G}_{1}(w_{1}))$
assuming $\tau<\delta$. Since $f_{1}$ is strongly convex in both $w_{1}$ and $w_{0}$ under Assumption 3, by the Prékopa–Leindler inequality [saumard2014log, Theorem 3.8], so is $\mathcal{G}_{1}$ with
$\nabla_{y}^{2}\mathcal{G}_{1}\succeq\frac{\int(\nabla_{y}^{2}f_{1}-\nabla_{yz}^{2}f_{1}(\nabla_{z}^{2}f_{1})^{-1}\nabla_{zy}^{2}f_{1})\exp(-f_{1}(y,z))dz}{\int\exp(-f_{1}(y,z))dz},$ (C.1)
@@
$=\frac{\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi})}{1+\tau(\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi}))}I$ (C.2)
@@
$c_{i}$ $=\frac{c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi})}{1+\tau(c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi}))},\quad i=1,2,3,\dots,N$ (C.3)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}.$ (C.4)
@@
$\frac{1}{2}V_{\pi}(w_{N})+\mathcal{G}_{N}(w_{N})\succeq\left(\frac{1}{2}\alpha_{\pi}+c_{N}\right)I.$ (C.5)
In order to understand the limit $\tau\rightarrow 0$ of the recursion (C.3), notice that it can be seen as a two step time discretisation, i.e.
$\tilde{c}_{i}$ $=c_{i-1}-\tau b^{2}$ (C.6)
$c_{i}$ $=\frac{\tilde{c}_{i}}{1+\tau\tilde{c}_{i}}$ (C.7)
with $b^{2}=\frac{|\alpha_{h}-L_{\pi}|}{2}$, recalling that $\alpha_{h}-L_{\pi}<0$. Also, (C.6) corresponds to an Euler discretisation of the ODE $\frac{dc_{t}}{dt}=-b^{2}$ with time step $\tau$. Notice also that (C.7) is the exact solution of $\frac{dc_{t}}{dt}=-c_{t}^{2}$ initialised at $\widehat{c}_{i}$ at time $t=\tau$ since if $c_{t}=c_{0}(1+tc_{0})^{-1}$ then differentiating with respect to $t$ yields $\dot{c}_{t}=-c_{0}^{2}(1+tc_{0})^{-2}=-c_{t}^{2}$. Therefore, the iteration (C.6)-(C.7) corresponds to an operator splitting based discretisation of the ODE
$\frac{dc_{t}}{dt}=-c_{t}^{2}-b^{2}.$ (C.8)
Notice that for the case $\frac{dc_{t}}{dt}=-c_{t}^{2}-b^{2}$, the ODE has no stable fixed point, and continues decreasing to -$\infty$ as $t\rightarrow\infty$, consistent with the findings of [Kolesnikov2001] that the W flow does not preserve strong log-concavity uniformly in time (aside from the Ornstein-Ulenbeck flow). We will then focus on characterising the time horizon over which strong convexity is preserved. Integrating both sides of the ODE and assuming $c_{t}^{2}+b^{2}\neq 0$ for all $t$ yields
$c_{t^{\ast}}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt^{\ast}\right).$
To summarise, the exact solution to (C.8) is given by
$c_{t}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt\right),$ (C.9)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}$ (C.10)
@@
$\widehat{\mu}_{t}$ $\propto\exp\left(-\frac{1}{2}V_{\pi}-\mathcal{G}_{t}\right)$
$=:\exp\left(-\mathcal{E}_{t}\right)$
@@
When $\alpha_{h}-L_{\pi}<0$, there exists some $t^{\ast}$ beyond which strong convexity111of the intermediate densities is lost, that is, $c_{t}\leq 0$ for all $t\geq t^{\ast}$. Due to the monotonocity of the tan function, $t^{\ast}$ is easily obtained as
$t^{\ast}$ $=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)-\frac{1}{b}\tan^{-1}\left(0\right)$ (C.11)
$=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)$ (C.12)
which is strictly positive since $b,c_{0}>0$.
@@
Proof.
Consider a fixed time interval $[0,T]$. We will characterise strong log-concavity for $t\in[0,T]$ under (2.1) and then demonstrate that this holds for all $T>0$.
Consider the sequence of densities due to the W-FR splitting scheme (2.6). Recall that $\widehat{\nu}_{i}$ is the density after a step of W flow and $\mu_{i}$ is the density after a step of FR flow. We first show that strong log-concavity is preserved for a sequence of steps with step size $\{\tau_{i}\}_{i=1,2,\dots}$, and then show that strong log-concavity is preserved as $\tau_{i}\rightarrow 0$. As will become clear in the remainder of the proof, such an iteration dependent step size is needed due to the fact that the time horizon over which the W flow preserves strong log-concavity depends on the relative convexity of the initial and target potentials.
@@
$\widehat{c}_{1}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-b\tau_{1}\right)$
@@
$\succeq\left(\left(\tfrac{1-\delta_{1}-e^{-\tau_{1}}}{2}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}\right)I$
$=:\alpha_{d,1}I$

### v2

@@ B.1 Moment ODEs and analytic solutions
(iii) We will demonstrate the reasoning for $\mathrm{KL}(\mu_{n\gamma}||\pi)$, noting that the same calculations hold for $\mathrm{KL}(\nu_{n}||\pi)$ and $\mathrm{KL}(\eta_{n}||\pi)$ due to Lemma 3.1 with $E_{n}$ replaced by $E_{n}^{\beta}$ and $E_{n}^{\alpha}$ respectively. Using the results of Lemma 3.1,
$\mathrm{KL}(\mu_{n\gamma}||\pi)$ $=\frac{1}{2}\left[-\log\det(I+C_{\pi}^{-1}E_{n})+Tr[C_{\pi}^{-1}E_{n}]+\varepsilon_{0}^{\top}E_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}E_{n}C_{\pi}^{-1}E_{n}e^{n\gamma C_{\pi}^{-1}}E_{0}^{-1}\varepsilon_{0}\right].$ (2.24)
As $C_{\pi}^{-1}$ is symmetric p.d, it is orthogonally diagonalisable, i.e. $C_{\pi}^{-1}=P^{\top}\Lambda P$ with $\lambda_{i}=\Lambda(i,i)$ denoting the $i$th eigenvalue of $C_{\pi}^{-1}$ and $0<\lambda_{1}<\lambda_{2}<\lambda_{3}\dots<\lambda_{d}$ and $p_{i}$ the $i$th column of $P$ where $P^{\top}P=I$. Now, consider the third term in the above, using the expression for $E_{n}$,
$\varepsilon_{0}^{\top}E_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}E_{n}C_{\pi}^{-1}E_{n}e^{n\gamma C_{\pi}^{-1}}E_{0}^{-1}\varepsilon_{0}$ $=e^{-2n\gamma}Tr[PU_{n}^{-1}E_{0}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}E_{0}^{-1}U_{n}^{-1}P^{\top}e^{-2n\gamma\Lambda}\Lambda]$
$=\sum_{i=1}^{d}\lambda_{i}e^{-2n\gamma(1+\lambda_{i})}N_{n}(i,i),$
@@
where $U_{n}$ is as defined in (i) and $N_{n}:=PU_{n}^{-1}E_{0}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}E_{0}^{-1}U_{n}^{-1}P^{\top}$. Furthermore, since $C_{\pi}^{-1}E_{n}=C_{\pi}^{-1}C_{n\gamma}-I$ has eigenvalues larger than 1-1, we have $\det(I+C_{\pi}^{-1}E_{n})=e^{Tr[\log(I+C_{\pi}^{-1}E_{n})]}$, and
$-\log\det(I+C_{\pi}^{-1}E_{n})=-Tr[\log(I+C_{\pi}^{-1}E_{n})]=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}Tr[(C_{\pi}^{-1}E_{n})^{k}]$
@@
whenever $C_{\pi}^{-1}E_{n}||<1$ (so that the series converges), which holds for sufficiently large $n$. Then the first and second terms in (2.24) are given by
$-\log\det(I+C_{\pi}^{-1}E_{n})+Tr[C_{\pi}^{-1}E_{n}]$ $=\sum_{k=2}^{\infty}\frac{(-1)^{k}}{k}Tr[(C_{\pi}^{-1}E_{n})^{k}]$
$=\sum_{k=2}^{\infty}\frac{(-1)^{k}}{k}e^{-n\gamma k}\,Tr[(D_{n}M_{n})^{k}],$
where $M_{n}:=PU_{n}^{-1}P^{\top}$ and $D_{n}:=e^{-n\gamma\Lambda}\Lambda e^{-n\gamma\Lambda}$ is a diagonal matrix. Starting with $k=2$,
$Tr[(D_{n}M_{n})^{2}]$ $=\sum_{i=1}^{d}\sum_{j=1}^{d}D_{n}(i,i)D_{n}(j,j)M_{n}(i,j)M_{n}(j,i)=\sum_{i=1}^{d}\sum_{j=1}^{d}\lambda_{i}\lambda_{j}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{n}(i,j)M_{n}(j,i),$
where the terms in the summand decay exponentially with $n$ at rate $2n\gamma(\lambda_{i}+\lambda_{j})$. A simple induction argument shows that as $k$ increases, the summands decay at increasing rates, since $\lambda_{i}>0$. Combining all and using the same reasoning for $\mathrm{KL}(\nu_{n}||\pi)$ yields that $\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ is given by
$\frac{\sum_{i=1}^{d}\sum_{j=1}^{d}\lambda_{i}\lambda_{j}e^{-2n\gamma(\lambda_{i}+\lambda_{j}+1)}M_{n}^{\beta}(i,j)M_{n}^{\beta}(j,i)+\sum_{i=1}^{d}\lambda_{i}e^{-2n\gamma(\lambda_{i}+1)}N_{n}^{\beta}(i,i)+T_{\beta}}{\sum_{i=1}^{d}\sum_{j=1}^{d}\lambda_{i}\lambda_{j}e^{-2n\gamma(\lambda_{i}+\lambda_{j}+1)}M_{n}(i,j)M_{n}(j,i)+\sum_{i=1}^{d}\lambda_{i}e^{-2n\gamma(\lambda_{i}+1)}N_{n}(i,i)+T}$
$=\frac{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+1)}\lambda_{i}N_{n}^{\beta}(i,i)+\mathcal{O}(e^{-2n\gamma(2\lambda_{1}+1)})}{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+1)}\lambda_{i}N_{n}(i,i)+\mathcal{O}(e^{-2n\gamma(2\lambda_{1}+1)})}$
$=\frac{\lambda_{1}N_{n}^{\beta}(1,1)+\mathcal{O}(e^{-2n\gamma\min(\lambda_{1},\lambda_{2}-\lambda_{1})})}{\lambda_{1}N_{n}(1,1)+\mathcal{O}(e^{-2n\gamma\min(\lambda_{1},\lambda_{2}-\lambda_{1})})}$
@@
where $T_{\beta}:=\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}e^{-n\gamma k}Tr[(D_{n}M_{n}^{\beta})^{k}]$ and likewise for $T$ with $M_{n}^{\beta}$ replaced by $M_{n}$ and $M_{n}^{\beta}$ and $N_{n}^{\beta}$ take the same form as $M_{n}$ and $N_{n}$ respectively but with $U_{n}$ replaced by $U_{n}^{\beta}$. Note that in the third line, we have divided top and bottom by $e^{-2n\gamma(\lambda_{1}+1)}$ and also noting that $U_{n}\rightarrow E_{0}^{-1}+\Omega C_{\pi}^{-1}$ and $U_{n}^{\beta}\rightarrow E_{0}^{-1}+\Omega^{\beta}C_{\pi}^{-1}$ as $n\rightarrow\infty$. Taking the limit of the above expression as $n\rightarrow\infty$ yields (3.12).
@@ Appendix C Preservation of log-concavity
Our proof strategy involves starting with a sequential splitting of the WFR flow and using Girsanov theorem to characterise the intermediate density due to the W flow. This is based on a similar proof strategy used to show that the filtering density satisfies a Poincaré inequality (see Lemma 5.1 in [PRS21]). In Appendix C.1, we show that when the initial density is “sufficiently” strongly log-concave (to be made precise), there exists a finite time interval over which the W flow maintains log-concavity. Although the W flow cannot be expected to preserve log-concavity uniformly in time (unless $\pi$ is Gaussian), the properties of the FR flow can be exploited to maintain log-concavity uniformly in time. The next lemma shows that the FR does exactly this, and then the proof in Appendix C.2 combines the results.
@@
The first step will be to use Girsanov theorem to characterise the law of the overdamped Langevin diffusion
$dY_{t}=-\nabla V_{\pi}(Y_{t})dt+\sqrt{2}d\tilde{W}_{t}$ (3.1)
where $\{\tilde{W}_{t}\}_{t\geq 0}$ is a standard Brownian motion. We will instead consider the time-rescaled Langevin diffusion with $s=\frac{t}{2}$,
$dX_{s}=\nabla U(X_{s})ds+dW_{s}.$ (3.2)
with $U=-\frac{1}{2}V_{\pi}$. Note that (3.2) has the same invariant density as (3.1). Let $\widehat{\mu}_{s}$ denote the law of $X_{s}$, so then by an application of Girsanov’s theorem (see [ØKS03, Exercise 8.15]),
$\widehat{\mu}_{s}$ $=\int_{C[0,s]}\exp\left(\int_{0}^{s}\nabla U(W_{u})dW_{u}-\frac{1}{2}\int_{0}^{s}|\nabla U(W_{u})|^{2}du\right)\mu_{W}(dw)$
$=\int_{C[0,s]}\exp\left((U(W_{s})-U(W_{0})-\frac{1}{2}\int_{0}^{s}(\Delta U+|\nabla U|^{2})(W_{u})du\right)\mu_{W}(dw)$
where $\mu_{W}(dw)$ denotes the Wiener measure on the space of continuous paths on $[0,s]$, $C[0,s]$. Note that Girsanov’s theorem applies here since Novikov’s condition is easily verified as $U$ satisfies a linear growth condition $|\nabla U|\leq c_{1}(1+|x|)$ for some $c_{1}>0$ under Assumption 2. In the second line, we have used Itô’s formula to obtain
$U(W_{s})-U(W_{0})=\int_{0}^{s}\frac{1}{2}\Delta U(W_{u})du+\int_{0}^{s}\nabla U(W_{u})dW_{u}.$
Consider now a discretisation of the time interval $[0,s]$ with step size $\tau$ and $N$ steps, $0=s_{0},s_{1},s_{2},\dots,s_{N-1},s_{N}=s$ with $s_{i}-s_{i-1}=\tau$ for all $i=1,2,3,\dots,N$. From now on, we use the notation $\widehat{\mu}_{i}$ to denote an approximation to $\widehat{\mu}_{s_{i}}$ (and likewise for other quantities). We arrive at the approximation
$\widehat{\mu}_{s}\approx$ $\int_{C[0,t]}\exp\left(\frac{1}{2}(V_{\pi}(w_{0})-V_{\pi}(w_{N}))\right)\prod_{i=1}^{N-1}\exp\left(-\frac{1}{2}R(w_{i})\right)\mu_{W}(dw)$
@@
$\mu_{i}(w_{i}):=\left\{\begin{array}[]{lr}\int Q(w_{{i-1}})\,q_{\tau}(w_{{i-1}},w_{i})\,\mu_{i-1}(w_{i-1})\,dw_{{i-1}},&i=2,3,\dots\\ \int Q(w_{0})\,q_{\tau}(w_{0},w_{i})\,\exp\left(\frac{1}{2}V_{\pi}(w_{0})\right)\,\exp(-V_{0}(w_{0}))\,dw_{0},&i=1\end{array}\right$.
@@
$R(w)$ $:=-\frac{1}{2}\Delta V_{\pi}(w)+\frac{1}{4}|\nabla V_{\pi}(w)|^{2}$
@@
$=\exp(-\mathcal{G}_{1}(w_{1})),$
assuming $\tau<\delta$. Since $f_{1}$ is strongly convex in both $w_{1}$ and $w_{0}$ under Assumption 3, by the Prékopa–Leindler inequality [SW14, Theorem 3.8], so is $\mathcal{G}_{1}$ with
$\nabla_{y}^{2}\mathcal{G}_{1}\succeq\frac{\int(\nabla_{y}^{2}f_{1}-\nabla_{yz}^{2}f_{1}(\nabla_{z}^{2}f_{1})^{-1}\nabla_{zy}^{2}f_{1})\exp(-f_{1}(y,z))dz}{\int\exp(-f_{1}(y,z))dz},$ (3.3)
@@
$=\frac{\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi})}{1+\tau(\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi}))}I$ (3.4)
@@
$c_{i}$ $=\frac{c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi})}{1+\tau(c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi}))},\quad i=1,2,3,\dots,N$ (3.5)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}.$ (3.6)
@@
$\frac{1}{2}V_{\pi}(w_{N})+\mathcal{G}_{N}(w_{N})\succeq\left(\frac{1}{2}\alpha_{\pi}+c_{N}\right)I.$ (3.7)
In order to understand the limit $\tau\rightarrow 0$ of the recursion (3.5), notice that it can be seen as a two step time discretisation, i.e.
$\tilde{c}_{i}$ $=c_{i-1}\pm\tau b^{2}$ (3.8)
$c_{i}$ $=\frac{\tilde{c}_{i}}{1+\tau\tilde{c}_{i}}$ (3.9)
with $b^{2}=\frac{|\alpha_{h}-L_{\pi}|}{2}$ and the plus or minus in (3.8) depending on the sign of $\alpha_{h}-L_{\pi}$. Also, (3.8) corresponds to an Euler discretisation of the ODE $\frac{dc_{s}}{ds}=\pm b^{2}$ with time step $\tau$. Notice also that (3.9) is the exact solution of $\frac{dc_{s}}{ds}=-c_{s}^{2}$ initialised at $\widehat{c}_{i}$ at time $s=\tau$ since if $c_{s}=c_{0}(1+tc_{0})^{-1}$ then differentiating with respect to $s$ yields $\dot{c}_{s}=-c_{0}^{2}(1+sc_{0})^{-2}=-c_{s}^{2}$. Therefore, the iteration (3.8)-(3.9) corresponds to an operator splitting based discretisation of the ODE
$\frac{dc_{s}}{ds}=-c_{s}^{2}\pm b^{2}.$ (3.10)
Notice that for the case $\frac{dc_{s}}{ds}=-c_{s}^{2}-b^{2}$, the ODE has no stable fixed point, and continues decreasing to -$\infty$ as $t\rightarrow\infty$. For this case, we focus on characterising the time horizon over which strong convexity is preserved. Integrating both sides of (3.10) yields
$c_{s}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bs\right),$ (3.11)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}$ (3.12)
@@
$\widehat{\mu}_{t}$ $\propto\exp\left(-\frac{1}{2}V_{\pi}-\mathcal{G}_{t}\right)=:\exp\left(-\mathcal{E}_{t}\right)$
@@
Now there exists some $s^{\ast}$ beyond which strong convexity of the intermediate densities is lost, that is, $c_{s}\leq 0$ for all $s\geq s^{\ast}$. Due to the monotonocity of the tan function, $s^{\ast}$ is easily obtained as
$s^{\ast}$ $=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)-\frac{1}{b}\tan^{-1}\left(0\right)=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)$ (3.13)
which is strictly positive since $b,c_{0}>0$. A change of time-scale back to $t=2s$ to obtain the concavity constants for (3.1) yields the result (4.1) and the time until this holds is given by $t^{\ast}=\frac{1}{2b}\tan^{-1}\left(\frac{c_{0}}{b}\right)$.
Now consider the case when $\alpha_{h}-L_{\pi}>0$, for which the corresponding ODE of interest is given by
$\frac{dc_{s}}{ds}=-c_{s}^{2}+b^{2}.$
A simple separation of variables calculation yields
$c_{s}$ $=b\left(\frac{K-e^{-2bs}}{K+e^{-2bs}}\right),\qquad K=\frac{b+c_{0}}{b-c_{0}}.$
Once again, applying the time-rescaling $t=2s$ yields (4.3) and $c_{t}>0$ for all $t$.
@@
We first consider the case $\alpha_{h}-L_{\pi}<0$. Consider the sequence of densities due to the W-FR splitting scheme (2.6). Recall that $\widehat{\nu}_{i}$ is the density after a step of W flow and $\mu_{i}$ is the density after a step of FR flow. We first show that strong log-concavity is preserved for a sequence of steps with step size $\{\tau_{i}\}_{i=1,2,\dots}$, and then show that strong log-concavity is preserved as $\tau_{i}\rightarrow 0$, over an infinite time horizon. As will become clear in the remainder of the proof, such an iteration dependent step size is needed due to the fact that the time horizon over which the W flow preserves strong log-concavity depends on the relative convexity of the initial and target potentials.
@@
$\widehat{c}_{1}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-2b\tau_{1}\right)$
@@
$\succeq\left(\left(\tfrac{1-\delta_{1}-e^{-\tau_{1}}}{2}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}\right)I=:\alpha_{d,1}I$